# Signatures of magnon hydrodynamics in an atomically-thin ferromagnet

### v1

@@ IV Discussion
In this work, we measure the dissipative magnetic response of atomically-thin CrCl based on the coherence of nearby NV centers in diamond. The central experimental finding is the observation of a continuous enhancement of magnetic noise with decreasing temperature that is only weakly suppressed by 30-fold increase in the externally applied magnetic field. The data are interpreted as a signature of exchange-collision induced hydrodynamic magnon transport, in which the NV center is employed as a magnon viscometer. We further corroborate this interpretation by discussing in the supplement how other mechanisms, including the damping of single magnon Sparks et al. (1961); Jermain et al. (2017), magnetic vortices Lu et al. (2020), magnetization tunneling Vitale et al. (1994), exotic magnetic particles Dusad et al. (2019), and spin glass physics Reim et al. (1986); Quilliam et al. (2008) all fail to fully capture the experimental phenomenology (see supplement XV.A). While we do not fully account for a potential coupling of the sample signal to the spin bath of the background (primarily P1 centers Bauch et al. (2020); Dwyer et al. (2022)), our analysis in supplement XV.A indicates that such a coupling is unlikely to be the origin of the anomalous noise trend. We also note that the microscopic calculation of the hydrodynamic transport coefficients of a strongly-interacting magnon gas presents a challenge to future theory.
This work provides a first glimpse of hydrodynamic magnons, and in particular evidence for viscous damping of a low-lying magnon sound mode. We also provide concrete predictions for the sound mode velocity, motivating future efforts towards its direct detection, as well as ultimately its potential application in magnon-based devices. Finally, the NV-based sensing protocol introduced here may be further extended to study magnetic dynamics in a wide range of materials.
@@ V Acknowledgments
The authors would like to thank I. Esterlis and J. Rodriguez-Nieva for fruitful discussions, and M. Tschudin for assistance with the experiment. This work is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). A. Y. is also partly supported by the Gordon and Betty Moore Foundation through Grant GBMF 9468, and by the U.S. Army Research Office (ARO) MURI project under grant number W911NF-21-2-0147. R. X. is also partly supported by the Army Research Office under Grant numbers: W911NF-22-1-0248. N.M. is supported by an appointment to the Intelligence Community Postdoctoral Research Fellowship Program at Harvard University administered by Oak Ridge Institute for Science and Education. The work of P.E.D. is sponsored by the Army Research Office and was accomplished under Grant Number W911NF-21-1-0184. A.M. and E.D. are supported by the SNSF project 200021212899 and the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number UeM019-1. FM acknowledges support from the NSF through a grant for ITAMP at Harvard University. PJH acknowledges support by the 2DMAGIC MURI FA9550-19-1-0390, the National Science Foundation (DMR-1809802), the STC Center for Integrated Quantum Materials (NSF grant no. DMR-1231319), the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant GBMF9463, and the Ramon Areces Foundation.
@@
R.X., N.M., and A.Y. conceived the experiment. R.X., N.M., and R.K. performed the experiment and data analysis. P.E.D., A.M., and F.M. developed the theoretical interpretation of the experiment, with critical input from A.Y., E.D., and M.D.L. L-Q.X. and R.X. fabricated the samples. D.R.K. and D.M. grew and characterized the CrCl crystals. K.W. and T.T. provided the hBN crystals. All authors contributed to the writing of the manuscript.

### v2

@@ IV Discussion
The presence of a hydrodynamic magnon sound mode in CrCl3 provides a consistent picture of several complementary observations in the present work. In particular, the anomalous trend in the temperature-dependent noise generated by monolayer CrCl3 indicates the presence of a viscously-damped hydrodynamic magnon sound mode (Fig. 3A,D). Our spectroscopy measurements on multilayer CrCl3 provide direct evidence for this mode (Fig. 5C).
The quantitative properties of a hydrodynamic magnon gas are set by two transport coefficients—the magnon sound velocity and viscosity. The present work reveals two exciting opportunities to explore these coefficients via both theoretical and experimental developments. Notably, the estimated experimental velocity of the sound mode (150 m/s) is slower than that expected in our semi-classical theory (300 m/s). Part of this discrepancy may reflect the crudeness of our experimental estimate, but it is also likely that the system exhibits physics beyond our present understanding. In particular, our prediction of the sound mode velocity is based on the dispersion of a bare magnon band. The presence of a hydrodynamic magnon sound mode, however, originates from strong interaction effects between the magnons, which are in turn expected to enhance the magnon effective mass Goldhirsch and Yakhot (1980) and slow down the sound mode relative to the results of our spin-wave theory.
Furthermore, experimentally extracting the other key transport coefficient, viscosity , from either the equilibrium noise measurements (Fig. 3) or the driven experiments (Fig. 5) requires precise knowledge of the magnon mass. Thus, a more complete theoretical picture of magnon renormalization effects and independent experimental measurement of the high-momentum magnon dispersion are critical for a quantitative understanding of the hydrodynamic magnon regime.
More broadly, this work serves as a starting point for exploring the physics of magnon sound in magnets Michel and Schwabl (1970); Reiter (1968); Rodriguez-Nieva et al. (2022), a platform with key advantages over the more familiar case of second-sound in superfluids Hohenberg and Martin (1965). For example, coherent magnon sound waves could be realized at ambient conditions in magnets with large exchange constants Buczek et al. (2020). Such longitudinal spin waves even exhibit properties that may find value in spintronic, magnonic, and caloritronic devices Ulloa et al. (2019); Sano and Matsuo (2023).
@@ V Acknowledgments
The authors would like to thank I. Esterlis, J. Rodriguez-Nieva, M. Ziffer, and B. Halperin for fruitful discussions, N. R. Reed for support on the forward Fourier method, and M. Tschudin for assistance with the experiment. This work is supported by the Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). A. Y. is also partly supported by the Gordon and Betty Moore Foundation through Grant GBMF 9468, and by the U.S. Army Research Office (ARO) MURI project under grant number W911NF-21-2-0147. R. X. is also partly supported by the Army Research Office under Grant numbers: W911NF-22-1-0248. N.M. is supported by an appointment to the Intelligence Community Postdoctoral Research Fellowship Program at Harvard University administered by Oak Ridge Institute for Science and Education. The work of P.E.D. is sponsored by the Army Research Office and was accomplished under Grant Number W911NF-21-1-0184. A.M. and E.D. are supported by the SNSF project 200021212899, the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number UeM019-1 and NCCR SPIN, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 225153). F.M. acknowledges support from the NSF through a grant for ITAMP at Harvard University. PJH acknowledges supports by by the Army Research Office MURI (grant no. W911NF2120147), the 2DMAGIC MURI (grant no. FA9550-19-1-0390), the National Science Foundation (grant no. DMR-1809802), the Office of Naval Research (grant no. N000142412440), the Ramón Areces Foundation and the Gordon and Betty Moore Foundation’s EPiQS Initiative through grant no. GBMF9463.
@@
R.X., N.M., and A.Y. conceived the experiment. R.X., N.M., and R.K. performed the experiment and data analysis. P.E.D., A.M., and F.M. developed the theoretical interpretation of the experiment, with critical input from A.Y., E.D., and M.D.L. L-Q.X. and R.X. fabricated the samples under the supervision of P.J.-H.. D.R.K. and D.M. grew and characterized the CrCl3 crystals under the supervision of P.J.-H.. K.W. and T.T. provided the hBN crystals. All authors contributed to the writing of the manuscript.